OpenStudy (anonymous):

Algebra 2 Help please. ((x+1)^4 -1)

OpenStudy (anonymous):

Your statement is incomplete. We have no instructions.

OpenStudy (anonymous):

Simplify...

OpenStudy (anonymous):

This is the difference of two squares ((x + 1)^2 + 1))((x + 1)^2 - 1) the second one factors down further ((x + 1)^2 + 1)((x + 1) + 1)((x + 1) - 1)

OpenStudy (anonymous):

((x + 1)^2 + 1) = (x + 1)(x + 1) + 1 x^2 + 2x + 1 + 1 (x^2 + 2x + 2)

OpenStudy (anonymous):

The other two are ((x + 1) + 1) = (x + 2) ((x + 1) - 1) = x

OpenStudy (anonymous):

So x(x + 2)(x^2 + 2x + 2)

OpenStudy (anonymous):

I got x^4+4x^3+6x^2+4x...but i dont knoww..

OpenStudy (anonymous):

I think it needs to be factored as I did above.

OpenStudy (anonymous):

I dont understand how you did it.

OpenStudy (anonymous):

These ones are harder to spot but it is basically a^2 - b^2. (x+1)^4 is basically the square of (x+1)^2 and 1 you know can be squared to equal 1